# Mining Glitch Tokens in Large Language Models via Gradient-based Discrete Optimization

### v2

@@ 4.1 Experimental Setup
Target LLM. We used a diverse set of LLMs from five different model families to evaluate the performance of our glitch token detection approach. The selected models include Meta’s Llama series llama2 ; llama3_1 , Alibaba’s Qwen models qwen2 ; qwen2_5 , Google’s Gemma models gemma2 , Microsoft’s Phi-3 models phi3 , and Mistral models mistral_7B ; mistral2024nemo . The details are presented in Table 1.
@@
Table 1: Target LLMs used in the experiments.
Evaluation Metrics. To assess the performance of our glitch token detection method, we employ the Top-N Precision metric, denoted as Precision@N. This metric evaluates the proportion of true glitch tokens within the top $N$ predictions made by the algorithm. It is calculated as:
@@
For example, Precision@1000 measures the precision of the top 1000 predicted tokens. This metric not only captures the accuracy of detecting true glitch tokens among high-ranking candidates but also reflects the effectiveness of a method under limited query budgets. As such, it is particularly well-suited for evaluating the performance of glitch token detection approaches, where efficiency and precision are crucial.
@@
4.2 Experimental Results
@@
Llama-3.1-8B Precision@1000 8.2% 50.4% 38.2%
Precision@2000 7.05% 42.35% 49.87%
Llama-2-7B Precision@1000 6.6% 9.5% 31.67%
Precision@2000 7.2% 9.15% 26.53%
Qwen2.5-7B Precision@1000 11.7% 0.4% 45.47%
Precision@2000 12.0% 0.2% 44.89%
Qwen2-7B Precision@1000 9.7% 0.4% 63.77%
Precision@2000 11.35% 0.2% 52.6%
Phi-3-mini-128k Precision@1000 2.0% 42.0% 31.53%
Precision@2000 2.25% 28.2% 21.15%
Phi-3.5-mini Precision@1000 3.0% 43.0% 28.0%
Precision@2000 3.75% 30.15% 25.11%
Gemma-2-2b-it Precision@1000 1.4% 71.9% 83.27%
Precision@2000 1.4% 55.5% 57.65%
Gemma-2-9b-it Precision@1000 4.0% 72.9% 90.17%
Precision@2000 3.1% 67.5% 70.57%
Mistral-7B-v0.3 Precision@1000 3.8% 6.6% 18.73%
Precision@2000 3.05% 7.45% 14.62%
Mistral-Nemo Precision@1000 4.5% 0.0% 57.03%
Precision@2000 4.45% 28.75% 61.48%
Average Precision@1000 5.49% 29.71% 48.78%
Precision@2000 5.56% 26.95% 42.45%
Table 2: Precision@1000 and Precision@2000 comparison of methods across different models. All results are averaged over three runs.
Table 2 presents a comprehensive comparison of GlitchMiner against state-of-the-art methods GlitchHunter and Magikarp across various LLM architectures. The results demonstrate the superior performance and versatility of our proposed approach.
GlitchMiner consistently outperforms both baselines in detecting glitch tokens, with particularly impressive results on newer model architectures. For instance, on the Gemma-2-9b model, GlitchMiner achieves a Precision@1000 of 90.17% and Precision@2000 of 70.57%, significantly surpassing both GlitchHunter and Magikarp. Similar performance gains are observed across the Qwen and Llama model families.
@@
The effectiveness of GlitchMiner is particularly evident in cases where Magikarp struggles, such as with the Qwen and Mistral models. This highlights a key advantage of our approach: its independence from manually defined priors. Unlike Magikarp, which relies heavily on assumed token distributions, GlitchMiner’s entropy-based loss function and local discrete optimization strategy allow it to adapt to the unique characteristics of each model’s vocabulary and token distribution.
This adaptability is crucial when dealing with the diverse landscape of modern LLMs, each with its own architectural nuances and token vocabularies. GlitchMiner’s consistent performance across different model families demonstrates its robustness and generalizability, addressing a significant limitation of previous approaches.
Remark. These experimental results not only showcase GlitchMiner’s superior precision in glitch token detection but also highlight its ability to generalize across diverse LLM architectures. This combination of accuracy and adaptability represents a significant advancement in the field of glitch token detection, providing a more reliable and versatile tool for improving the robustness and security of LLMs.
@@ 4.3 Ablation Study
Effect of Neighborhood Size. In figure 4, we can see that as the neighborhood size $K$ increases, there is a gradual decline in precision. However, if $K$ is too small, the exploration is insufficient, which also reduces accuracy, as shown in the experiments with Llama-3.1-8B. Therefore, balancing $K$ is crucial to ensure both broad enough exploration and sufficient approximation accuracy.
@@
Effect of Batch Size. As shown in figure 5, different models exhibit varying sensitivity to batch size $B$. For example, Llama-3.1-8B and Qwen2.5-7B perform best at $B=4$, while Phi-3-mini peaks at $B=8$. Gemma-2-2B, on the other hand, reaches its highest performance at $B=16$ and maintains strong results at larger $B$ values. This indicates that the optimal $B$ value depends on the specific model architecture.
@@ 5 Conclusion
In conclusion, we introduce GlitchMiner, a novel method for efficiently mining glitch tokens in large language models through gradient-based discrete optimization. By leveraging entropy as the loss function and integrating a local search strategy, GlitchMiner accurately identifies tokens that deviate from normal distributions without relying on manually defined priors. Extensive experimental results demonstrate that GlitchMiner outperforms existing state-of-the-art methods across various model architectures, showcasing higher precision and adaptability. Ablation studies further confirm the contributions of key components to the overall performance. The consistent results across different models emphasize the robustness and generalizability of GlitchMiner. Future work may explore applying this method to other high-dimensional tasks that require efficient discrete optimization.

### v3

@@ 4.1 Experimental Setup
Test LLMs. We used a diverse set of LLMs from five different model families to evaluate the performance of our glitch token detection approach. The selected models include Meta’s Llama series llama2 ; llama3_1 , Alibaba’s Qwen models qwen2 ; qwen2_5 , Google’s Gemma models gemma2 , Microsoft’s Phi-3 models phi3 , and Mistral models mistral_7B ; mistral2024nemo . The details are presented in Table 1.
@@
Table 1: Test LLMs used in the experiments.
Evaluation Metrics. To assess the performance of our glitch token detection method, we employ the Detected@N metric. This metric measures the absolute number of true glitch tokens identified within the top $N$ predictions made by the algorithm. It is calculated as:
@@
For example, Detected@1000 measures the number of true glitch tokens discovered within the top 1000 predicted tokens. This metric directly reflects a method’s effectiveness in identifying glitch tokens under practical constraints, as it captures both the detection capability and query efficiency of the approach. By comparing the number of glitch tokens detected under the same query budget ($N$), we can effectively evaluate different methods’ ability to maximize token discovery while minimizing computational resources. This makes it an ideal metric for real-world applications where efficient detection of glitch tokens is crucial.
@@
Initialization Strategy in Experiments. To ensure stable and consistent comparisons across runs, we initialize the search with the token exhibiting the smallest $\ell_{2}$ norm in the embedding space, based on prior observations that such tokens often exhibit glitch-like behaviors. However, as shown in Figure 4, we found that GlitchMiner remains robust to different initialization choices, achieving similar performance even with random starting points.
4.2 Main Results
@@
Llama-3.1-8B-Instruct Detected@1000 25 664 568
Detected@2000 56 935 1164
Llama-2-7B-chat-hf Detected@1000 61 100 319
Detected@2000 126 186 532
Qwen2.5-7B-Instruct Detected@1000 75 1000 1000
Detected@2000 180 1893 1839
Qwen2-7B-Instruct Detected@1000 96 999 1000
Detected@2000 191 1842 1847
Gemma-2-2b-it Detected@1000 23 678 744
Detected@2000 35 984 1019
Gemma-2-9b-it Detected@1000 29 623 775
Detected@2000 45 983 1089
Phi-3.5-mini-instruct Detected@1000 20 393 396
Detected@2000 44 496 516
Phi-3-mini-128k-instruct Detected@1000 26 398 404
Detected@2000 55 489 517
Mistral-7B-Instruct-v0.3 Detected@1000 6 110 219
Detected@2000 19 130 302
Mistral-Nemo-Instruct-2407 Detected@1000 48 574 695
Detected@2000 79 918 976
Average Detected@1000 40.9 553.9 612.0
Detected@2000 93.0 885.6 980.1
Table 2: Detected@1000 and Detected@2000 comparison of methods across different models.
Table 2 displays the performance comparison of GlitchMiner with the state-of-the-art methods GlitchHunter and Magikarp across different LLM architectures using the Detected@N metric. The results highlight GlitchMiner’s strong and versatile detection capabilities.
In the majority of cases, GlitchMiner achieved the highest Detected@2000 score, consistently outperforming both baselines in terms of identified glitch tokens within the top 2000 predictions. This is particularly evident in models such as Llama-2-7B-chat-hf, where GlitchMiner achieved a Detected@2000 of 532, surpassing Magikarp by a notable margin. For the Qwen and other models, GlitchMiner maintained robust performance, demonstrating adaptability across different LLM architectures.
@@
These experimental findings underscore GlitchMiner’s ability not only to accurately detect glitch tokens but also to generalize across a diverse set of model architectures. This adaptability and precision position GlitchMiner as a powerful tool for enhancing the robustness and security of LLMs.
@@ 4.3 Ablation Study
Effect of Neighborhood Size. We analyzed the impact of neighborhood size $K$ on detection performance. As shown in Figure 2, increasing $K$ generally leads to a decline in Detected@1000 values across models. This trend indicates that as $K$ grows, the Taylor approximation becomes less effective, resulting in reduced prediction accuracy.
@@
Figure 4: Effect of Initialization Method on GlitchMiner’s Detected@1000 score.
Effect of Batch Size. As shown in Figure 3, the performance of GlitchMiner remains relatively stable as batch size $B$ increases. Notably, even with $B=1$, GlitchMiner achieves effective detection results, indicating that it can make accurate predictions without relying on a large batch size.
Effect of Initialization Token. As shown in Figure 4, GlitchMiner’s performance remains stable across different initialization tokens. The red dots represent the minimum $\ell_{2}$ norm initialization, while the orange dots show three random trials. For most models, random initialization results are close to the minimum $\ell_{2}$ norm, indicating that GlitchMiner achieves consistent detection accuracy regardless of the initialization approach.
@@ 5 Conclusion
In this paper, we introduced GlitchMiner, a novel framework for detecting glitch tokens in LLMs through gradient-based discrete optimization. Our method combines entropy-based loss functions with a local search strategy to navigate token space efficiently, improving both detection accuracy and computational efficiency. Experimental results across diverse LLMs highlight GlitchMiner’s robustness and versatility compared to existing methods. Our findings confirm GlitchMiner’s effectiveness in identifying glitch tokens across architectures, demonstrating its potential for enhancing the security and reliability of LLM-based applications. Future work could explore extending this approach to other high-dimensional discrete optimization tasks.
Broader Impact
The potential impact of this research extends beyond mere academic interest. By improving the detection of glitch tokens, GlitchMiner contributes to enhancing the overall reliability and safety of LLMs in real-world applications. This is particularly crucial in high-stakes domains such as healthcare, finance, and autonomous systems, where the consequences of model errors can be severe. Furthermore, our approach opens new avenues for understanding and mitigating vulnerabilities in AI systems, potentially influencing the development of more robust and trustworthy AI technologies in the future.
Acknowledgments
We would like to express our sincere gratitude to Sander Land for his valuable insights and feedback on this work, particularly regarding the evaluation metrics. His suggestions were instrumental in refining our approach and improving the clarity of our results. We also wish to thank Guanlin Li for his valuable suggestions on our manuscript. His insights were crucial for improving the content and structure of the paper.